Find the largest integer such that 4 more than 3 times the integer is less than 93.
Found 2 solutions by palanisamy, Theo:
Answer by palanisamy(496)

Let x be the required number
From the given data, we get
3x+4<93
3x<93-4
3x<89
x<89/3
x=29
Therefore the required number is 29

Answer by Theo(13342)

3*x + 4 < 93

subtract 4 from both sides to get:

3*x < 89

divide both sides by 3 to get:

x < 89/3

89/3 = 29.67

x = 29 = largest integer less than 29.67

3 * 29 + 4 = 87 + 4 = 91 < 93

3 * 30 + 4 = 90 + 4 = 94 NOT < 93

number is 29.